Quiz: 3.2 Quiz Time Remaining: 01:09:39 Submit Qui This Question: 1 pt 4 of 7 This Quiz: 7 pts possib of the population of a large country is math phobic. If two people are randomly selected, what is the probability both are math phobic? What is the probability at least one is math phobic? (a) The probability that both will be math phobic is Round to four decimal places as needed) (b) The probability that at least one person is math phobic is Round to four decimal places as needed) . -s* | 1.1 Ja :/İ : .. (..) More 2 Enter your answer in each of the answer boxes 10:49 PM
Answer #1

a)P(both will be math phobic) =0.09*0.09=0.0081

b)P(at least one will be math phobic)=1-P(none is math phobic)=1-0.91*0.91=0.1719